OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-2,2) and (-1,10)

OpenStudy (anonymous):

use the form y-y1=m(x-x1) find the slope: 8 (y-2)=8(x+2) and when we simplify we have y=8x+18